OpenStudy (anonymous):

Write the standard form of the equation of the circle with endpoints of a diameter at the points of (5,8) and (-9,11)

OpenStudy (amistre64):

add the points and divide by 2 to find the center

OpenStudy (amistre64):

( 5, 8 ) +(-9,11) ---------- (-4,19); /2 to get (-2, 19/2) alot of it is just now: \((x+2)^2+(y-(19/2))^2 = r^2\)

OpenStudy (amistre64):

to get r; subtract the center from an endpoint ( 5,8 ) -(-2,19/2) ( 5,16/2) -(-2,19/2) ----------- 7, -3/2 ; now we squre them and add them (7)^2 + (-3/2)^2 = r^2

OpenStudy (anonymous):

I'm getting (x+2)\[(x+2)^{2}+ (y-19/2)^{2}=205/2\]

OpenStudy (anonymous):

but not sure that is right

OpenStudy (amistre64):

it should come out to 51' 1/4 ; which equals 205/4

OpenStudy (anonymous):

ugh, so close

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

thanks